I have included screenshots of my work. I was hoping someone could check my answers, and assist me if they are wrong, and help me thrrough the ones I dont have answered. Thanks.Patterson and Associates You have been hired by Patterson & Associates, a small events planning company that recently had a fDepreciation 1. Determine which depreciation method is shown in each schedule on the Patterson & Associates panel. Then matchFinal Questions Review the depreciation schedules on the Patterson & Associates panel, then answer the following questions. 1
Answer #1

Answer.

Schedule A B C
Useful Life 5 4 7
Residual Value $2,000 $3,375
Assets Cost $20,000 $54,000 $47,250
Operating Hours 10,500

Explanation: In depreciation Schedule A, total depreciation expense is $18,000, Useful Life of asset is 5 years and in initial years depreciation expenses are more as compared to the following years. Thus, it represents Double Declining Balance method of depreciation which is followed for assets generating greater revenues in earlier years as compared to the following years.

Depreciation Rate = (1 \div 5) \times 2 = 0.4 or 40%

First Year Depreciation = $,8,000, Book Value of Assets (cost) =?

Depreciation Expense for year 1 = Book Value \times Depreciation rate

$8,000 = Book Value \times 0.4 OR

Book Value(cost) = $8,000 \div 0.4 = $20,000

Residual Value = Book Value (Cost) - Total Depreciation

= 20,000 - 18,000 = $2,000

Schedule B: The useful life of Asset is only 4 years and total depreciation expense for four years is $50,625 (sum of first four year Dep.). The total cost of asset is $54,000. Thus, Residual Value is Cost less Depreciation. That is $54,000 - $50,625 = $3,375

Depreciation Schedule C: Operating Hours

The useful life of Asset is only 4 years and depreciation rate per hour is $4.5. This asset will depreciate over 7 years, and has no salvage value. Thus, we can calculate Operating hours by dividing Total depreciation expense by depreciation rate per hour.

Operating Hours = Total Depreciation / Rate per hour = 10,500 Hours

Final Quest.

The total depreciation would be $50625 and Depreciation for third year $13500.